# Infinity categories with duality and hermitian multiplicative infinite loop space machines

### v3

@@ 7 Dualities on the $\infty$-category of spectra
In the following we prove that $\pi_{0}(\mathrm{pic}(S)^{BC_{2}})$ contains ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2},$ where ${\mathbb{Z}}_{2}$ is the ring of 2-adic numbers. By Lin’s Theorem [19, Theorem 1.1] there is an isomorphism $\pi_{0}(S^{BC_{2}})\cong{\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$ and we prove that the canonical map of spectra $S\to\mathrm{pic}(S)$ that sends the unit to $S[1]$ induces an injection (Proposition 7.7):
@@
Let ${\mathcal{D}}$ be a stable $\infty$-category that admits small limits and $\mathrm{X}\in{\mathcal{D}}.$ Choosing a base point in the space $BC_{2}$, the retraction *$**\to BC_{2}\to*$ of spaces induces a retraction $\mathrm{X}\to\mathrm{X}^{BC_{2}}\to\mathrm{X}$ in ${\mathcal{D}}.$ Thus the cofiber sequence of pointed spaces (*$(*)_{+}\to(BC_{2})_{+}\to BC_{2}$ induces via the cotensoring of ${\mathcal{S}}_{*}$ on ${\mathcal{D}}$ a split exact sequence $\mathrm{X}^{BC_{2}}_{*}\to\mathrm{X}^{BC_{2}}\to\mathrm{X}$, where the first object is the pointed cotensor. In particular, the sequence
@@
By [19, Theorem 1.1] there is an isomorphism $\pi_{0}(S^{BC_{2}})\cong{\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$. The split fiber sequence
@@
The functor $\nu:\mathrm{Sp}^{C_{2}}\to\mathrm{Sp}[C_{2}]$ that forgets along the functor $BC_{2}\subset\mathsf{Fin}[C_{2}]\to\mathrm{Span}(\mathsf{Fin}[C_{2}])$ admits a fully faithul right adjoint $\iota:\mathrm{Sp}[C_{2}]\to\mathrm{Sp}^{C_{2}}$ [24, Proposition 6.17.] such that $b:=\iota\circ\nu$, which we call Borel-completion, is $\mathrm{Sp}^{C_{2}}$-linear.

### v4

@@ 7 Dualities on the $\infty$-category of spectra
In the following we prove that $\pi_{0}(\mathrm{pic}(S)^{BC_{2}})$ contains ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2},$ where ${\mathbb{Z}}_{2}$ is the ring of 2-adic numbers. By Lin’s Theorem [23, Theorem 1.1] there is an isomorphism $\pi_{0}(S^{BC_{2}})\cong{\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$ and we prove that the canonical map of spectra $S\to\mathrm{pic}(S)$ that sends the unit to $S[1]$ induces an injection (Proposition 7.7):
@@
Let ${\mathcal{D}}$ be a stable $\infty$-category that admits small limits and $\mathrm{X}\in{\mathcal{D}}.$ Choosing a base point in the space $BC_{2}$, the retraction $*\to BC_{2}\to*$ of spaces induces a retraction $\mathrm{X}\to\mathrm{X}^{BC_{2}}\to\mathrm{X}$ in ${\mathcal{D}}.$ Thus the cofiber sequence of pointed spaces $(*)_{+}\to(BC_{2})_{+}\to BC_{2}$ induces via the cotensoring of ${\mathcal{S}}_{*}$ on ${\mathcal{D}}$ a split exact sequence $\mathrm{X}^{BC_{2}}_{*}\to\mathrm{X}^{BC_{2}}\to\mathrm{X}$, where the first object is the pointed cotensor. In particular, the sequence
@@
By [23, Theorem 1.1] there is an isomorphism $\pi_{0}(S^{BC_{2}})\cong{\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$. The split fiber sequence
@@
The functor $\nu:\mathrm{Sp}^{C_{2}}\to\mathrm{Sp}[C_{2}]$ that forgets along the functor $BC_{2}\subset\mathsf{Fin}[C_{2}]\to\mathrm{Span}(\mathsf{Fin}[C_{2}])$ admits a fully faithul right adjoint $\iota:\mathrm{Sp}[C_{2}]\to\mathrm{Sp}^{C_{2}}$ [28, Proposition 6.17.] such that $b:=\iota\circ\nu$, which we call Borel-completion, is $\mathrm{Sp}^{C_{2}}$-linear.